# Unveiling Hallucination in Text, Image, Video, and Audio Foundation Models: A Comprehensive Review

### v1

@@ Abstract
The rapid advancement of foundation models (FMs) across language, image, audio, and video domains has shown remarkable capabilities in diverse tasks. However, the proliferation of FMs brings forth a critical challenge: the potential to generate hallucinated outputs, particularly in high-stakes applications. The tendency of foundation models to produce hallucinated content arguably represents the biggest hindrance to their widespread adoption in real-world scenarios, especially in domains where reliability and accuracy are paramount. This survey paper presents a comprehensive overview of recent developments that aim to identify and mitigate the problem of hallucination in FMs, spanning text, image, video, and audio modalities. By synthesizing recent advancements in detecting and mitigating hallucination across various modalities, the paper aims to provide valuable insights for researchers, developers, and practitioners. Essentially, it establishes a clear framework encompassing definition, taxonomy, and detection strategies for addressing hallucination in multimodal foundation models, laying the foundation for future research in this pivotal area.
@@ 1.1 Hallucination
Hallucination refers to instances where FMs generate content that appears plausible and emulates human-like patterns but lacks a coherent understanding of the underlying context or factual grounding Xu et al. (2024b). These hallucinated outputs can manifest in various forms, ranging from trivial factual inaccuracies to more severe cases where the model generates entirely imaginary or absurd content. This issue is not limited to textual modalities but extends across diverse domains, including images, videos, and audio generated by large-scale foundation models. The root causes of hallucination are multifaceted, potentially stemming from biases in training data, limited access to current information, or the model’s inherent constraints in comprehending and generating contextually precise responses. Deploying these powerful models without addressing their hallucination tendencies can have severe consequences, perpetuating misinformation, leading to incorrect conclusions, and potentially causing adverse effects in critical applications. Addressing hallucinations in FMs has become an active research area and a key focus for development efforts, with ongoing efforts to mitigate this behavior. Strategies employed include fine-tuning models on domain-specific data, leveraging diverse training data to enhance model robustness, and developing improved evaluation metrics to identify and reduce hallucination tendencies.
@@ 1.2 Types of Hallucination
Hallucinations in large foundation models can manifest in various forms, each posing unique challenges and requiring tailored strategies for mitigation. These types of hallucinations can be categorized as follows: Contextual disconnection Zhang et al. (2023d) describes a situation in which the output or content produced by a model across different modalities is inconsistent or out of sync with the context that the user or the input data have provided or expected. Semantic distortion  Tjio et al. (2022) refers to a type of inconsistency or error in generated content where the semantics or underlying meaning of the input is misrepresented or altered in the output. Content hallucination Moernaut et al. (2018)is the term used to describe a phenomenon seen in generative models when features or elements that are generated as output are either unreal given the context or absent from the input data. Factual inaccuracy  Zhang et al. (2023d) describes a kind of error seen in generative models when information that is inaccurate, deceptive, or at odds with the known facts appears in the generated output. Figure 1 illustrates various types of hallucinations.
Figure 1: Types of hallucinations illustrated: This image shows several types of hallucinations. Proper explanations of hallucinations are indicated as hallucinated elements (HE) and are highlighted in bold red text.
@@ 1.3 Motivation and Contributions
Most of the existing survey papers have explored hallucination in the context of large language models (LLMs) Huang et al. (2023)Tonmoy et al. (2024). Recent studies have shown that hallucination can also occur in vision, audio, and video foundation models, highlighting the need for a comprehensive understanding of this challenge across multiple modalities Liu et al. (2024a)Sahoo et al. (2024)Rawte et al. (2023b). To address this gap, our survey aims to provide a holistic and multimodal perspective on the hallucination challenge in FMs. This review comprehensively examines existing research across language, vision, video, and audio domains to understand the mechanisms, detection methods, and mitigation strategies for hallucination in FMs. It serves as a vital resource for researchers and practitioners, aiding in the development of more robust AI solutions. Additionally, it includes a detailed taxonomy diagram in Figure. 2 and a summarized Table 1 illustrating recent advancements across different modalities. The contributions of this survey paper are as follows:
@@
  • Identify the key factors and mechanisms that contribute to the emergence of hallucination across different modalities.
@@
  • We have provided a comprehensive summary of the methodologies pertaining to hallucination techniques in large foundational models in Table 1, detailing their approaches to hallucination detection, mitigation, task considerations, datasets utilized, and evaluation metrics employed. This will offer readers a concise overview of recent advancements in this field.
The paper is structured as follows: Section 2 delves into hallucinations in LLMs, while Section 3 explores large vision language models. Additionally, Section 4 and Section 5 examine hallucinations in large video and large audio models, respectively. Section 6 presents a discourse on the ramifications of hallucinations, exploring whether they pose advantages or disadvantages. Subsequently, Section 7 addresses the limitations of the current work, while Section 8 outlines potential future directions. Finally, Section 9 presents the concluding remarks.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style= rotate=90, child anchor=north, parent anchor=south, anchor=center , , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Hallucinations, for tree=fill=bg16, text width=14em [ Text §2, for tree=fill=red!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ FACTOID Rawte et al. (2024b), for tree=fill=pg56,text width=28em ] [ FACTOR Muhlgay et al. (2023), for tree=fill=pg56,text width=28em ] [ FactCHD Chen et al. (2023b), for tree=fill=pg56,text width=28em ] [ HalluQA Cheng et al. (2023), for tree=fill=pg56,text width=28em ] [ HaluEval Li et al. (2023b), for tree=fill=pg56,text width=28em ] [ SelfCheckGPT Manakul et al. (2023), for tree=fill=pg56,text width=28em ] [ FACTSCORE Min et al. (2023), for tree=fill=pg56,text width=28em ] [ FacTool Chern et al. (2023), for tree=fill=pg56,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ DoLa Chuang et al. (2023), for tree=fill=bg4,text width=28em ] [ CoK Li et al. (2023d), for tree=fill=bg4,text width=28em ] [ MixAlign Knowledge , for tree=fill=bg4,text width=28em ] [ SELF-FAMILIARITY Luo et al. (2023), for tree=fill=bg4,text width=28em ] [ CoVe Dhuliawala et al. (2023), for tree=fill=bg4,text width=28em ] [ HALOCHECK Elaraby et al. (2023), for tree=fill=bg4,text width=28em ] [ Instructional Prompting Varshney et al. (2023), for tree=fill=bg4,text width=28em ] [ PURR Chen et al. (2023a), for tree=fill=bg4,text width=28em ] [ LLM-AUGMENTER Peng et al. (2023), for tree=fill=bg4,text width=28em ] ] ] [ Image §3, for tree=fill=orange!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ HallusionBench Guan et al. (2023), for tree=fill=dandyshade,text width=28em ] [ GAVIE Liu et al. (2023), for tree=fill=dandyshade,text width=28em ] [ PhD Liu et al. (2024b), for tree=fill=dandyshade,text width=28em ] [ VHTest Huang et al. (2024), for tree=fill=dandyshade,text width=28em ] [ M-HalDetect Gunjal et al. (2024), for tree=fill=dandyshade,text width=28em ] [ ChartBench Xu et al. (2023b), for tree=fill=dandyshade,text width=28em ] [ MSG-MCQ Lu et al. (2024), for tree=fill=dandyshade,text width=28em ] [ MMVP Tong et al. (2024), for tree=fill=dandyshade,text width=28em ] [ REVO-LION Liao et al. (2023), for tree=fill=dandyshade,text width=28em ] [ CIEM Hu et al. (2023), for tree=fill=dandyshade,text width=28em ] [ FAITHSCORE Jing et al. (2023), for tree=fill=dandyshade,text width=28em ] [ POPE Li et al. (2023f), for tree=fill=dandyshade,text width=28em ] [ VQA Changpinyo et al. (2022), for tree=fill=dandyshade,text width=28em ] [ HaELM Wang et al. (2023), for tree=fill=dandyshade,text width=28em ] [ NOPE Lovenia et al. (2023), for tree=fill=dandyshade,text width=28em ] [ MMHAL-BENCH Sun et al. (2023), for tree=fill=dandyshade,text width=28em ] [ TouchStone Bai et al. (2023), for tree=fill=dandyshade,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ HalluciDoctor Yu et al. (2023a), for tree=fill=bg8,text width=28em ] [ LURE Zhou et al. (2023), for tree=fill=bg8,text width=28em ] [ MARINE Zhao et al. (2024), for tree=fill=bg8,text width=28em ] [ FDPO Gunjal et al. (2024), for tree=fill=bg8,text width=28em ] [ ViGoR Yan et al. (2024), for tree=fill=bg8,text width=28em ] [ HA-DPO Zhao et al. (2023), for tree=fill=bg8,text width=28em ] [ VIGC Wang et al. (2024), for tree=fill=bg8,text width=28em ] [ Data Centric Approach Lu et al. (2024), for tree=fill=bg8,text width=28em ] [ MoF Tong et al. (2024), for tree=fill=bg8,text width=28em ] [ InternLM-XComposer Zhang et al. (2023b), for tree=fill=bg8,text width=28em ] [ VCD Leng et al. (2023), for tree=fill=bg8,text width=28em ] [ Factually Augmented RLHF Sun et al. (2023), for tree=fill=bg8,text width=28em ] [ ObjMLM Dai et al. (2022), for tree=fill=bg8,text width=28em ] ] ] [ Video §4 , for tree=fill=bg5, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ EMScore Shi et al. (2022), for tree=fill=bg7,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ FactVC Liu and Wan (2023), for tree=fill=bg13,text width=28em ] [ CLearViD Chuang and Fazli (2023), for tree=fill=bg13,text width=28em ] [ MGAT He et al. (2022), for tree=fill=bg13,text width=28em ] ] ] [ Audio §5 , for tree=fill=lime!50, text width=18em [ Detection, for tree=fill=bg1, text width=13.5em [ PAM Deshmukh et al. (2024), for tree=fill=bg15,text width=28em ] [ CompA Ghosh et al. (2023), for tree=fill=bg15,text width=28em ] ] [ Mitigation, for tree=fill=bg2, text width=13.5em [ MusicLDM Chen et al. (2024), for tree=fill=bg14,text width=28em ] [ Re-AudioLDM Yuan et al. (2024), for tree=fill=bg14,text width=28em ] [ SECap Xu et al. (2024a), for tree=fill=bg14,text width=28em ] [ RECAP Ghosh et al. (2024), for tree=fill=bg14,text width=28em ] [ Cacophony Zhu and Duan (2024), for tree=fill=bg14,text width=28em ] [ EnCLAP Kim et al. (2024), for tree=fill=bg14,text width=28em ] ] ] ]
@@ 2 Hallucination in Large Language Models
Despite the progress of LLMs, a notable challenge persists in their proneness to hallucinate, impeding their practical implementation. For instance, the illustration in Figure 3 exemplifies the generated response by the LLM, showcasing indications of hallucination.
@@ 2.1 Hallucination Detection and Mitigation
Identifying hallucinations in LLMs is crucial for ensuring the credibility and reliability of their results, especially in scenarios requiring factual correctness. Existing fact-checking methods often rely on complex modules or external databases, requiring either output probability distributions or interfacing with external sources. SelfCheckGPT Manakul et al. (2023) offers a zero-resource black-box solution for detecting hallucinations in any LLM without relying on external resources. This method operates on the principle that an LLM familiar with a topic will produce consistent and comparable facts in its responses. In contrast, randomly sampled responses from an unfamiliar topic are likely to contain contradicting and hallucinated facts. Continuing the exploration of methods for passage-level hallucination detection, Yang et al. (2023) proposed a novel self-check approach based on reverse validation, aiming to identify factual errors without external resources automatically. They introduced a benchmark, Passage-level Hallucination Detection (PHD), generated using ChatGPT and annotated by human experts to assess different methods. Assessing the accuracy of long text generated by LLMs is challenging because it often contains both accurate and inaccurate information, making simple quality judgments insufficient. To address this, Min et al. (2023) introduced FACTSCORE (Factual Precision in Atomicity Score), a novel evaluation method that breaks down text into individual facts and measures their reliability. Huang and Chang (2023) introduced a unique strategy to mitigate hallucination risks in LLMs by drawing parallels with established web systems. They identified the absence of a "citation" mechanism, which refers to acknowledging or referencing sources or evidence, as a significant gap.
@@
Addressing the need to identify factual inaccuracies in LLM-generated content, Rawte et al. (2024b) develop a multi-task learning (MTL) framework, integrating advanced long text embeddings like e5-mistral-7b-instruct, along with models such as GPT-3, SpanBERT, and RoFormer. This MTL approach exhibited a substantial performance gain, achieving an average accuracy improvement of 40% on the FACTOID benchmark when compared to leading textual entailment methods. Hallucination mitigation efforts have predominantly relied on empirical methods, leaving uncertainty regarding the possibility of complete elimination. To tackle this challenge, Xu et al. (2024b) introduces a formal framework defining hallucination as discrepancies observed between computable LLMs and a ground truth function. Through this framework, the study examines existing hallucination mitigation strategies and their practical implications for real-world LLM deployment. Rawte et al. (2024c) introduced the Sorry, Come Again (SCA) prompting technique to address hallucination in contemporary LLMs. SCA enhances comprehension by employing optimal paraphrasing and injecting [PAUSE] tokens to delay LLM generation. It analyzes linguistic nuances in prompts and their impact on the hallucinated generation, highlighting the difficulties posed by prompts characterized by lower readability, formality, or concreteness. Rawte et al. (2023a) investigate how LLMs respond to factually correct and incorrect prompts, categorizing their hallucinations into mild, moderate, and alarming subcategories. Additionally, the paper introduced the Hallucination eLiciTation dataset, comprising 75,000 text snippets annotated by humans, and introduced a novel Hallucination Vulnerability Index metric.
2.2 Domain Specific Works
Hallucinations pose severe risks in critical fields such as healthcare, finance, and law. Reliability and accuracy are crucial in these sectors, as any form of hallucination can lead to significant and adverse consequences.
2.2.1 Healthcare
In response to the hallucinations in the medical domain LLMs, Pal et al. (2023) introduced the Medical Domain Hallucination Test (Med-HALT), a specialized benchmark dataset aimed at assessing and mitigating hallucinations. Med-HALT consists of a varied multinational dataset drawn from medical records spanning numerous countries, encompassing a total of seven datasets. Ahmad et al. (2023) outlined essential steps for creating dependable, trustworthy, and unbiased models, emphasizing the need for quantifying, validating, and mitigating hallucinations within the healthcare context. Ji et al. (2023) introduced an interactive self-reflection approach aimed at improving the accuracy and coherence of answers generated by medical question-answering systems using LLMs. Through knowledge acquisition and feedback on answer generation, this methodology enhances the factuality, consistency, and logical progression of responses.
2.2.2 Finance
An empirical investigation explored the inclination of LLMs to generate hallucinations while engaging in financial activities.  Kang and Liu (2023) conducted an empirical investigation into the hallucination tendencies of LLMs in financial tasks. Their study assessed LLMs’ proficiency in explaining financial concepts, querying historical stock prices, and examined the efficacy of methods like few-shot learning and prompt-based tool learning in mitigating hallucination.  Roychowdhury et al. (2023) proposed a novel Langchain-based approach aimed at transforming data tables into hierarchical textual data chunks, facilitating versatile financial question answering. The framework involves classifying user queries by intention, retrieving relevant data chunks, generating customized LLM prompts, and evaluating responses for hallucinations and confidence.
2.2.3 Legal
The conventional approach to abstractive text summarization typically employs an encoder-decoder architecture, wherein the encoder encapsulates the essence of the source text while the decoder generates the summary. However, this method may produce summaries containing irrelevant or inaccurate information, which poses a significant concern in legal contexts where accuracy is crucial. To address these issues, Feijo and Moreira (2023) introduced LegalSumm, which creates distinct "views" of the source text, trains summarization models to produce independent summaries, and employs an entailment module to assess their fidelity to the source. Deroy et al. (2023) investigated the readiness of LLMs for generating abstractive summaries of case judgments by applying SOTA models to Indian court cases. While abstractive models generally scored slightly higher, the authors noted inconsistencies and hallucinations in the generated summaries. Understanding the meaning of open-ended legal terms is very important for legal professionals. They often look at how such terms were used and interpreted in previous court cases.  Savelka et al. (2023) evaluate GPT-4’s performance in generating factually accurate, clear, and relevant explanations of legal terms in legislation. A comparison is made between a baseline approach, where GPT-4 directly explains a legal term, and an augmented approach employing a legal information retrieval module to provide contextual sentences from case law.  Dahl et al. (2024) presents the initial evidence on the frequency and types of inaccuracies in the legal domain, providing valuable insights for evaluating LLMs in legal contexts. By examining the structured format of American case law, the study assesses three major LLMs: GPT-3.5, PaLM 2, and Llama.

### v4

# A Comprehensive Survey of Hallucination in Large Language, Image, Video and Audio Foundation Models

@@ Abstract
The rapid advancement of foundation models (FMs) across language, image, audio, and video domains has shown remarkable capabilities in diverse tasks. However, the proliferation of FMs brings forth a critical challenge: the potential to generate hallucinated outputs, particularly in high-stakes applications. The tendency of foundation models to produce hallucinated content arguably represents the biggest hindrance to their widespread adoption in real-world scenarios, especially in domains where reliability and accuracy are paramount. This survey paper presents a comprehensive overview of recent developments that aim to identify and mitigate the problem of hallucination in FMs, spanning text, image, video, and audio modalities. By synthesizing recent advancements in detecting and mitigating hallucination across various modalities, the paper aims to provide valuable insights for researchers, developers, and practitioners. Essentially, it establishes a clear framework encompassing definition, taxonomy, and detection strategies for addressing hallucination in multimodal foundation models, laying the foundation for future research and development in this pivotal area.
@@ 1.1 Hallucination
Hallucination refers to instances where FMs generate plausible but factually incoherent or absurd content lacking proper context understanding Xu et al. (2024b). These hallucinated outputs can range from minor inaccuracies to completely imaginary content, manifesting across text, images, videos, and audio generated by large models. Root causes may include biases in training data, limited access to up-to-date information, or inherent model constraints in contextual comprehension and response generation. Deploying such powerful models without addressing hallucinations risks perpetuating misinformation, incorrect conclusions, and adverse effects in critical applications. Mitigating hallucinations has become an active research focus, with strategies like fine-tuning with domain-specific data, using diverse, robust training data, and developing improved evaluation metrics to identify and reduce hallucination tendencies.
@@ 1.2 Types of Hallucination
Hallucinations in large FMs can manifest in various forms and can be categorized as follows: Contextual disconnection: Zhang et al. (2023d) described a situation in which the output or content produced by a model across different modalities is inconsistent or out of sync with the context that the user or the input data provided or expected. Semantic distortion:  Tjio et al. (2022) refers to a type of inconsistency or error in generated content where the semantics or underlying meaning of the input is misrepresented or altered in the output. Content hallucination is the term used to describe a phenomenon seen in generative models when features or elements that are generated as output are either unreal given the context or absent from the input data Moernaut et al. (2018). Factual inaccuracy:  Zhang et al. (2023d) described a kind of error seen in generative models when information that is inaccurate, deceptive, or at odds with the known facts appears in the generated output. Figure 1 illustrates various types of hallucinations with examples.
Figure 1: Illustration of Hallucination types. Proper explanations of hallucinations are indicated as hallucinated elements (HE) and are highlighted in bold red text.
@@ 1.3 Motivation and Contributions
Most of the existing survey papers have explored hallucination in the context of large language models (LLMs) Huang et al. (2023)Chen et al. (2024b)Tonmoy et al. (2024). Recent studies have shown that hallucination can also occur in vision, audio, and video foundation models, highlighting the need for a comprehensive understanding of this challenge across multiple modalities Liu et al. (2024a)Sahoo et al. (2024b)Rawte et al. (2023). To address this gap, the present survey aims to provide a holistic and multimodal perspective on the hallucination challenge in FMs across language, vision, video, and audio domains. It serves as a vital resource for researchers and practitioners, aiding in the development of more robust AI solutions. Additionally, it includes a detailed taxonomy diagram in Fig. 2 and a summarized Table 1 (refer to section 9.1 of the appendix) illustrating recent advancements across different modalities. The contributions of this survey paper are as follows:
  • This survey provides a comprehensive overview of hallucination detection and mitigation techniques designed specifically for multimodal foundation models, extending beyond the typical focus solely on language models.
@@
  • Highlight the open challenges and future research directions in this critical area.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=15em,font=,, where level=2text width=14em,font=,, where level=3minimum width=10em,font=,, where level=4text width=26em,font=,, where level=5text width=20em,font=,, [ Hallucinations, for tree=fill=paired-light-red!80, text width=14em [ Text §2, for tree=fill=red!50, text width=18em [ Detection
[ FACTOR Muhlgay et al. (2023), for tree=fill=pg56,text width=28em ] [ FactCHD Chen et al. (2023b), for tree=fill=pg56,text width=28em ] ] [ Inference Classifier [ HalluQA Cheng et al. (2023), for tree=fill=pg56,text width=28em ] [ HaluEval Li et al. (2023b), for tree=fill=pg56,text width=28em ] ] [ Self-Evaluation [ SelfCheckGPT Manakul et al. (2023), for tree=fill=pg56,text width=28em ] ] [ Evidence Retrieval [ FACTSCORE Min et al. (2023), for tree=fill=pg56,text width=28em ] [ FacTool Chern et al. (2023), for tree=fill=pg56,text width=28em ] ] ] [ Mitigation
[ LLM-AUGMENTER Peng et al. (2023), for tree=fill=pg58,text width=28em ] ] [ Chain-of-Thought and Reasoning Enhancement [ CoK Li et al. (2023e), for tree=fill=pg58,text width=28em ] [ CoVe Dhuliawala et al. (2023), for tree=fill=pg58,text width=28em ] ] [ Alignment and Fine-Tuning [ MixAlign Knowledge , for tree=fill=pg58,text width=28em ] [ PURR Chen et al. (2023a), for tree=fill=pg58,text width=28em ] ] [ Self-Verification and Consistency Checking [ SELF-FAMILIARITY Luo et al. (2023), for tree=fill=pg58,text width=28em ] [ HALOCHECK Elaraby et al. (2023), for tree=fill=pg58,text width=28em ] ] [ Prompt Engineering and Instructional Techniques [ Instructional Prompting Varshney et al. (2023), for tree=fill=pg58,text width=28em ] ] ] ] [ Image §3, for tree=fill=orange!50, text width=18em [ Detection
[ PhD Liu et al. (2024b), for tree=fill=Dandelion,text width=28em ] [ REVO-LION Liao et al. (2023), for tree=fill=Dandelion,text width=28em ] ] [ Visual Question Answering
[ VQA Changpinyo et al. (2022), for tree=fill=Dandelion,text width=28em ] ] [ Benchmark and Evaluation Datasets
[ ChartBench Xu et al. (2023b), for tree=fill=Dandelion,text width=28em ] [ MMHAL-BENCH Sun et al. (2023), for tree=fill=Dandelion,text width=28em ] [ TouchStone Bai et al. (2023), for tree=fill=Dandelion,text width=28em ] ] [ Multimodal Evaluation
[ M-HalDetect Gunjal et al. (2024), for tree=fill=Dandelion,text width=28em ] ] [ Quantitative Metrics and Scoring Systems
[ POPE Li et al. (2023g), for tree=fill=Dandelion,text width=28em ] [ NOPE Lovenia et al. (2023), for tree=fill=Dandelion,text width=28em ] ] [ CIEM Hu et al. (2023), for tree=fill=Dandelion,text width=28em ] [ MSG-MCQ Lu et al. (2024), for tree=fill=Dandelion,text width=28em ] [ HaELM Wang et al. (2023), for tree=fill=Dandelion,text width=28em ] ] [ Mitigation
[ LURE Zhou et al. (2023), for tree=fill=lightgray,text width=28em ] [ VCD Leng et al. (2023), for tree=fill=lightgray,text width=28em ] ] [ Fine-Tuning and Post-Processing
[ HA-DPO Zhao et al. (2023), for tree=fill=lightgray,text width=28em ] [ InternLM-XComposer Zhang et al. (2023b), for tree=fill=lightgray,text width=28em ] ] [ Generative Adversarial Networks
[ MARINE Zhao et al. (2024), for tree=fill=lightgray,text width=28em ] ] [ Data-Centric Approaches
[ Data Centric Approach Lu et al. (2024), for tree=fill=lightgray,text width=28em ] ] [ Multi-Modal Fusion and Object-Level Modeling
[ ObjMLM Dai et al. (2022), for tree=fill=lightgray,text width=28em ] ] [ Reinforcement Learning
] ] ] [ Video §4 , for tree=fill=bg25, text width=18em [ Detection
] [ Mitigation
[ CLearViD Chuang and Fazli (2023), for tree=fill=pg60,text width=28em ] [ MGAT He et al. (2022), for tree=fill=pg60,text width=28em ] ] ] [ Audio §5 , for tree=fill=lime!50, text width=18em [ Detection
[ CompA Ghosh et al. (2023), for tree=fill=pg70,text width=28em ] ] [ Mitigation
[ Re-AudioLDM Yuan et al. (2024), for tree=fill=red!50,text width=28em ] ] [ Retrieval-Based Approaches
] [ Self-Supervised and Contrastive Learning
[ EnCLAP Kim et al. (2024), for tree=fill=red!50,text width=28em ] ] [ Cacophony Zhu and Duan (2024), for tree=fill=red!50,text width=28em ] ] ] ]
@@ 2 Hallucination in Large Language Models
Despite the progress of LLMs, a notable challenge persists in their proneness to hallucinate, impeding their practical implementation. For instance, the illustration in Figure 3 exemplifies the generated response by the LLM, showing indications of hallucination.
@@ 2.1 Hallucination Detection and Mitigation
Detecting hallucinations in LLMs is crucial for ensuring the credibility and reliability of their results, especially in scenarios requiring factual correctness. Existing fact-checking methods often rely on complex modules or external databases, requiring either output probability distributions or interfacing with external sources. The SelfCheckGPT method Manakul et al. (2023) offers a zero-resource black-box solution for detecting hallucinations in any LLM without relying on external resources. This method operates on the principle that an LLM familiar with a topic will produce consistent and comparable facts in its responses. In contrast, randomly sampled responses from an unfamiliar topic are likely to contain contradicting and hallucinated facts. Continuing the exploration of methods for passage-level hallucination detection, Yang et al. (2023) proposed a novel self-check approach based on reverse validation, aiming to automatically identify factual errors without external resources. They introduced a benchmark, Passage-level Hallucination Detection (PHD), generated using ChatGPT and annotated by human experts to assess different methods. Assessing the accuracy of long text generated by LLMs is challenging because it often contains both accurate and inaccurate information, making simple quality judgments insufficient. To address this, Min et al. (2023) introduced FACTSCORE (Factual Precision in Atomicity Score), a new evaluation method that breaks down text into individual facts and measures their reliability.  Huang and Chang (2023) introduced a unique strategy to mitigate hallucination risks in LLMs by drawing parallels with established web systems. They identified the absence of a "citation" mechanism in LLMs, which refers to acknowledging or referencing sources or evidence, as a significant gap.
@@
Addressing the need to identify factual inaccuracies in LLM-generated content, Rawte et al. (2024b) developed a multi-task learning (MTL) framework, integrating advanced long text embeddings like e5-mistral-7b-instruct, along with models such as GPT-3, SpanBERT, and RoFormer. This MTL approach demonstrated a 40% average improvement in accuracy on the FACTOID benchmark compared to leading textual entailment methods. Hallucination mitigation efforts have predominantly relied on empirical methods, leaving uncertainty regarding the possibility of complete elimination. To tackle this challenge, Xu et al. (2024b) introduced a formal framework defining hallucination as inconsistencies between computable LLMs and a ground truth function. The study examines existing hallucination mitigation strategies and their practical implications for real-world LLM deployment through this framework. Rawte et al. (2024c) introduced the Sorry, Come Again (SCA) prompting technique to address hallucination in contemporary LLMs. SCA enhances comprehension through optimal paraphrasing and injecting [PAUSE] tokens to delay LLM generation. It analyzes linguistic nuances in prompts and their impact on the hallucinated generation, emphasizing how prompts with lower readability, formality, or concreteness pose challenges.
Benchmark Evaluation: In certain instances, LLMs engage in a phenomenon termed "hallucination snowballing," where they fabricate false claims to rationalize prior hallucinations despite acknowledging their inaccuracy Li et al. (2023c). To empirically explore this phenomenon, Zhang et al. (2023a) devised three question-answering datasets spanning diverse domains, wherein ChatGPT and GPT-4 often furnish inaccurate answers alongside explanations featuring at least one false claim. Significantly, the study suggests that the language model can discern these false claims as incorrect. Another benchmark dataset, FactCHD Chen et al. (2023b), was introduced to detect fact-conflicting hallucinations within intricate inferential contexts. It encompasses a range of datasets capturing different factuality patterns and integrates fact-based evidence chains to improve assessment accuracy.
Figure 4: Four IVL-Hallu examples in Prompted Hallucination Dataset(PhD)  Liu et al. (2024b) including visuals and the matching question-answer pairs and hallucination elements (HE). While words annotated in red do not exist or do not match within the image, words annotated in green have correspondences within the image. Question, Answer, and Statement are denoted by the letters Q, A, and S, respectively.
3 Hallucination in Large Vision-Language Models
Large Vision-Language Models (LVLMs) have garnered significant attention in the AI community for their ability to handle visual and textual data simultaneously Ghosh et al. (2024c)Ghosh et al. (2024b)Sahoo et al. (2024a)Ghosh et al. (2024d)Ghosh et al. (2024a). Nonetheless, similar to LLMs, LVLMs also confront the issue of hallucination. Figure 4 illustrates an example of visual hallucination.
3.1 Hallucination Detection and Mitigation
Dai et al. (2022) investigated the issue of object hallucinations in Vision-Language Pre-training (VLP) models, where textual descriptions generated by these models contain non-existent or inaccurate objects based on input images.  Li et al. (2023g) revealed widespread and severe object hallucination issues and suggested that visual instructions may influence hallucination, noting that objects frequently appearing in visual instructions or co-occurring with image objects are more likely to be hallucinated. To enhance the evaluation of object hallucination, they introduced a polling-based query method called POPE, which demonstrates improved stability and flexibility in assessing object hallucination. The absence of a standardized metric for assessing object hallucination has hindered progress in understanding and addressing this issue. To address this gap, Lovenia et al. (2023) introduced NOPE (Negative Object Presence Evaluation), a novel benchmark for evaluating object hallucination in vision-language (VL) models through visual question answering (VQA). Utilizing LLMs, the study generates 29.5k synthetic negative pronoun (NegP) data for NOPE. It extensively evaluates the performance of 10 VL models in discerning the absence of objects in visual questions, alongside their standard performance on visual questions across nine other VQA datasets. Most existing efforts focused primarily on object hallucination, overlooking the diverse types of LVLM hallucinations. Liu et al. (2024b) delved into Intrinsic Vision-Language Hallucination (IVL-Hallu) and proposed several novel IVL-Hallu tasks categorized into four types: attribute, object, multi-modal conflicting, and counter-common-sense hallucination. To assess and explore IVL-Hallu, they introduced a challenging benchmark dataset and conducted experiments on five LVLMs, revealing their incapacity to effectively address the proposed IVL-Hallu tasks. To mitigate object hallucination in LVLMs without resorting to costly training or API reliance, Zhao et al. (2024) introduced MARINE, which is both training-free and API-free. MARINE enhances the visual understanding of LVLMs by integrating existing open-source vision models and utilizing guidance without classifiers to integrate object grounding features, thereby improving the precision of the generated outputs. Evaluations across six LVLMs reveal MARINE’s effectiveness in reducing hallucinations and enhancing output detail, validated through assessments using GPT-4V. Deng et al. (2024) introduced a CLIP-Guided Decoding (CGD) training-free approach to reduce object hallucination at decoding time.
HalluciDoctor Yu et al. (2023a) tackled hallucinations in Multi-modal Large Language Models (MLLMs) by using human error detection to identify and eliminate various types of hallucinations. By rebalancing data distribution via counterfactual visual instruction expansion, they successfully mitigate 44.6% of hallucinations while maintaining competitive performance. Despite proficiency in visual semantic comprehension and meme humor, MLLMs struggle with chart analysis and understanding. Addressing this, Xu et al. (2023b) proposed ChartBench, a benchmark assessing chart comprehension. ChartBench exposes MLLMs’ limited reasoning with complex charts, prompting the need for novel evaluation metrics like Acc+ and a handcrafted prompt, ChartCoT. Zhang et al. (2023b) introduced InternLM-XComposer, an LVLM aimed at designed to address the challenge of hallucination in image-text comprehension and composition. The performance of InternLM-XComposer’s text-image composition is evaluated through a robust procedure involving both human assessment and comparison to GPT4-Vision, with the model demonstrating competitive performance against solutions like GPT4-V and GPT3.5. Wang et al. (2024b) proposed the Instruction Contrastive Decoding (ICD) method to reduce hallucinations during LVLM inference. A recent study Huang et al. (2024a) proposed a novel decoding approach that introduces an over-confidence penalty and a retrospective allocation strategy to mitigate hallucination issues without requiring additional data or retraining.